# Improving LLM Reasoning through Interpretable Role-Playing Steering

### v1

@@ Abstract
Anyi Wang1, Dong Shu2, Yifan Wang1, Yunpu Ma1,3, Mengnan Du4 1Center for Information and Language Processing, LMU Munich 2Northwestern University 3Munich Center for Machine Learning (MCML) 4New Jersey Institute of Technology {anyi.wang, yifan.wang}@campus.lmu.de,dongshu2024@u.northwestern.edu, cognitive.yunpu@gmail.com,mengnan.du@njit.edu
@@ 1 Introduction
To address the limitations of prompt-based role-playing, we propose a stable and interpretable steering approach, Sparse Autoencoder Role-Playing Steering (SRPS). We use a pretrained Sparse Autoencoder (SAE) to analyze the internal activations of an LLM during role-play prompting. We identify the top-$k$ features that are most strongly and frequently activated by role-play prompts, and use them to construct a steering vector that captures the behavioral shift induced by role-playing. At inference time, this vector is added to the model’s residual stream with a tunable scaling factor, enabling stable steering of the model’s internal state toward the desired role-based reasoning behavior. To better understand why the top-$k$ features influence the model’s reasoning process, we use Neuronpedia Lin (2023) to interpret their semantic functions, offering insight into how role-playing enhances model performance. Finally, we examine accuracy variation across different sets of $k$ features, demonstrating the effectiveness of our feature ranking method. Extensive experiments show that our steering mechanism consistently enhances the model’s reasoning ability and achieves performance comparable to prompt-based role-playing in most cases. Our contributions can be summarized as follows:
@@ 2.1 Sparse Autoencoders
Steering methods have emerged as effective approaches for controlling and interpreting the behavior of LLMs Wu et al. (2024); Templeton et al. (2024); Arditi et al. (2024); Singh et al. (2024); Bi et al. (2024). SAEs Shu et al. (2025) have further strengthened this direction by offering a powerful tool for uncovering interpretable structures within LLMs. Cunningham et al. (2023) demonstrate that SAEs can extract highly interpretable and monosemantic features from LLM activations, enabling fine-grained analysis of model behavior. Building upon this, Bricken et al. (2023) apply dictionary learning techniques to decompose LLM representations into semantically meaningful components, facilitating a deeper understanding of their internal mechanisms. To enhance the reconstruction fidelity of SAEs without compromising interpretability, Rajamanoharan et al. (2024) introduce JumpReLU SAEs, which utilize a discontinuous activation function to achieve state-of-the-art performance on large models like Gemma2-9B Team et al. (2024). Furthermore, Gao et al. (2024) explore the scalability of SAEs, training models with up to 16 million latents on GPT-4 Achiam et al. (2023) activations and establishing clear scaling laws that balance sparsity and reconstruction accuracy. These advancements highlight the effectiveness of SAEs in interpreting and analyzing the complex behaviors of LLMs.
@@ 3.1 Extracting Role-Play Activations
We construct $N$ sample pairs from the training set for every dataset, each consisting of a positive and a negative sample. The positive sample includes a question prepended with a role-play prompt, while the negative sample is the same question without the role-play instruction. To avoid overfitting to the surface form of a single instruction, where the model memorizes its literal meaning instead of learning a generalizable behavior vector He et al. (2025), we design five semantically diverse prompt variants for each role.
@@ 3.2 Role-Relevant Feature Selection
$s_{i}=\mu_{i}+\beta\cdot\delta_{i},$ (3)
@@ 3.3 Steering Vector Construction
As sentence-level instructions contain rich and complex meanings spread across multiple features, using a single vector is often insufficient for effective steering He et al. (2025). Therefore, we use a set of vectors to better capture the full effect of role-playing. We define the overall steering shift as a weighted combination of the top-$k$ steering vectors, where the weight for each selected feature corresponds to its mean activation $\alpha_{i}$ across $N$ positive samples. The final steering shift vector $\mathbf{s}\in\mathbb{R}^{h}$ is computed as:
@@ 4.1 Experimental Settings
Baselines. We compare the performance of the steered model against two baselines: the model with the original question and the model using a prepended role-play prompt. Evaluations are conducted under three settings: few-shot CoT (4-shot), one-shot CoT, and zero-shot CoT. We append “Let’s think step by step” to the original question to elicit CoT reasoning in LLMs. For one-shot and few-shot exemplars, we follow Wei et al. (2022) to construct them (see Table 6 and 7 in Appendix A.1). For the prompt-based baseline, we prepend each of the five role-specific prompts designed for steering to the input question respectively, and select the best result among them for comparison. The model accuracy for each role-play prompt, along with the average and standard deviation across the five prompts, is reported in Table 8 in Appendix B.
Implementation Details. The number of input sample pairs $N$ is set to 1,000. We use pretrained SAEs from Gemma Scope Lieberum et al. (2024) and Llama Scope He et al. (2024) for latent activation extraction. We use the SAE with 131K dimension from layer 25 for Llama3.1-8B, 65K dimension from layer 25 for Gemma2-2B, and 131K dimension from layer 35 for Gemma2-9B. In our experiments, we choose the top 15 SAE latent features with the highest sensitivity scores for steering, as He et al. (2025) demonstrate that the model achieves the optimal performance when steering is performed using 15 latent dimensions. Other hyperparameter choices for all the experimental settings are summarized in Table 4 in Appendix C, where we further discuss the influence of different hyperparameters on the steering effect.
@@ 4.2 Benchmark Performance Comparison
Effectiveness of the Proposed SRPS. As shown in Table 1, our steering approach demonstrates superior performance, outperforming the original prompting in 25 out of 27 evaluation settings and achieving comparable results in the remaining two settings on the CSQA dataset. Compared to the prompt-based role-playing baseline, our method achieves better results in 17 out of 27 evaluation settings. These results highlight the effectiveness of our feature-based steering mechanism across models and tasks.
@@ 4.3 SAE Feature Analysis
To understand how the selected SAE features enhance model reasoning under role-playing, we conduct a semantic analysis of their activations using Neuronpedia Lin (2023). Specifically, we analyze features from Layer 35 of Gemma2-9B for arithmetic reasoning, and features from Layer 25 of Gemma2-2B for commonsense reasoning. Figure 3 presents the top-1 ranked feature across all evaluation settings in each domain, showing the corresponding positive and negative output logits. We found the extracted features align well with the expected domain knowledge for each reasoning task, validating the interpretability and task relevance of our steering mechanism. Figure 2 shows the impact of our steering method on model outputs. See Appendix D for more examples.
@@ 5 Conclusion
In this work, we have introduced an efficient and interpretable steering framework that leverages features extracted from a pretrained SAE to guide LLMs toward role-specific behavior. By constructing steering vectors from highly relevant features and injecting them into the model’s residual stream, our method enables fine-grained control over the model’s internal state without additional training. Extensive experiments across multiple models and reasoning benchmarks demonstrate that our approach consistently enhances model performance, yielding results on par with prompt-based role-playing. In addition, we analyze the semantic meanings of the selected features and uncover why role-playing enhances reasoning abilities in LLMs. Lastly, we evaluate steering performance using $k$ features selected from different ranking positions, and observe a declining trend in accuracy gains as lower-ranked features are used, validating the effectiveness of our feature selection method.
@@ A.3 Answer Extraction
Appendix B Role-play Prompting Baseline
@@
Appendix C Hyperparameters for Steering
We report all the parameters used for steering across different evaluation settings on the three benchmarks. As the number of features $k$ selected for steering is fixed at 15 in all experiments, we focus on comparing the remaining hyperparameters ($\theta$, $\beta$, and $\lambda$) in Table 4. Specifically, $\theta$ is the activation threshold used to determine whether a feature is considered active; $\beta$ is a weighting coefficient that balances the contribution of activation frequency and strength in feature selection; and $\lambda$ is a scaling factor that controls the strength of the steering vector when injected into the model’s residual stream.
Among the three models, Llama3.1-8B is the most sensitive to activation perturbation, and small changes in parameter values can lead to significant differences in steering performance. In contrast, the Gemma-2 family models, particularly Gemma2-9B, is less sensitive to activation modification. Among the three hyperparameters, $\beta$ and $\lambda$ have the greatest impact on steering performance. For models that are less sensitive to activation intervention, such as Gemma2-9B, larger values of $\beta$ and $\lambda$ are typically required to achieve optimal results. In general, $\theta$ is set within the range of 0 to 0.3 to maintain the effectiveness of steering vectors.

### v2

@@ Abstract
Anyi Wang1, Dong Shu2, Yifan Wang1, Yunpu Ma1,3, Mengnan Du4 1Center for Information and Language Processing, LMU Munich, 2Northwestern University 3Munich Center for Machine Learning (MCML) 4New Jersey Institute of Technology {anyi.wang, yifan.wang}@campus.lmu.de, dongshu2024@u.northwestern.edu, cognitive.yunpu@gmail.com, mengnan.du@njit.edu Corresponding authors
@@ 1 Introduction
To address the limitations of prompt-based role-playing, we propose a stable and interpretable steering approach, Sparse Autoencoder Role-Playing Steering (SRPS). We use a pretrained Sparse Autoencoder (SAE) to analyze the internal activations of an LLM during role-play prompting. We identify the top‑$k$ features that are most strongly and frequently activated by role-play prompts, and use them to construct a steering vector that captures the behavioral shift induced by role-playing. At inference time, this vector is added to the model’s residual stream with a tunable scaling factor, enabling stable steering of the model’s internal state toward the desired role-based reasoning behavior. To better understand why the top‑$k$ features influence the model’s reasoning process, we use Neuronpedia Lin (2023) to interpret their semantic functions, offering insight into how role-playing enhances model performance. Finally, we examine accuracy variation across different sets of $k$ features, demonstrating the effectiveness of our feature ranking method. Extensive experiments show that our steering mechanism consistently enhances the model’s reasoning ability and achieves performance comparable to prompt-based role-playing in most cases. Our contributions can be summarized as follows:
@@ 2.1 Sparse Autoencoders
Steering methods have emerged as effective approaches for controlling and interpreting the behavior of LLMs Wu et al. (2024); Templeton et al. (2024); Arditi et al. (2024); Singh et al. (2024); Bi et al. (2024); Li et al. (2025); He et al. (2025a). SAEs Shu et al. (2025) have further strengthened this direction by offering a powerful tool for uncovering interpretable structures within LLMs. Cunningham et al. (2023) demonstrate that SAEs can extract highly interpretable and monosemantic features from LLM activations, enabling fine-grained analysis of model behavior. Building upon this, Bricken et al. (2023) apply dictionary learning techniques to decompose LLM representations into semantically meaningful components, facilitating a deeper understanding of their internal mechanisms. To enhance the reconstruction fidelity of SAEs without compromising interpretability, Rajamanoharan et al. (2024) introduce JumpReLU SAEs, which utilize a discontinuous activation function to achieve state-of-the-art performance on large models like Gemma2-9B Team et al. (2024). Furthermore, Gao et al. (2024) explore the scalability of SAEs, training models with up to 16 million latents on GPT-4 Achiam et al. (2023) activations and establishing clear scaling laws that balance sparsity and reconstruction accuracy. These advancements highlight the effectiveness of SAEs in interpreting and analyzing the complex behaviors of LLMs.
@@ 3.1 Extracting Role-Play Activations
We construct $N$ sample pairs from the training set for every dataset, each consisting of a positive and a negative sample. The positive sample includes a question prepended with a role-play prompt, while the negative sample is the same question without the role-play instruction. To avoid overfitting to the surface form of a single instruction, where the model memorizes its literal meaning instead of learning a generalizable behavior vector He et al. (2025b), we design five semantically diverse prompt variants for each role.
@@ 3.2 Role-Relevant Feature Selection
$I_{i}=\mu_{i}+\beta\cdot\delta_{i},$ (3)
@@ 3.3 Steering Vector Construction
As sentence-level instructions contain rich and complex meanings spread across multiple features, using a single vector is often insufficient for effective steering He et al. (2025b). Therefore, we use a set of vectors to better capture the full effect of role-playing. We define the overall steering shift as a weighted combination of the top-$k$ steering vectors, where the weight for each selected feature corresponds to its mean activation $\alpha_{i}$ across $N$ positive samples. The final steering shift vector $\mathbf{s}\in\mathbb{R}^{h}$ is computed as:
@@ 4.1 Experimental Settings
Baselines. We compare the performance of our proposed steering method against two baselines: the model prompted with the original question, and the model using a prepended role-play prompt. Evaluations are conducted under three settings: few-shot CoT (4-shot), one-shot CoT, and zero-shot CoT. We append “Let’s think step by step” to the original question to elicit CoT reasoning in LLMs. For one-shot and few-shot exemplars, we follow Wei et al. (2022) to construct them (see Table 6 and 7 in Appendix A.1). For the prompt-based baseline, we prepend each of the five role-specific prompts designed for steering to the input question respectively, and select the best result among them for comparison. The model accuracy for each role-play prompt, along with the average and standard deviation across the five prompts, is reported in Table 8 in Appendix B.
Implementation Details. The number of input sample pairs $N$ is set to 1,000. We use pretrained SAEs from Gemma Scope Lieberum et al. (2024) and Llama Scope He et al. (2024) for latent activation extraction. Since deeper layers in LLMs integrate broader contextual information and capture higher-level representations, and SAEs with larger dimensions provide better interpretability of extracted features, we select the later layers and the highest available SAE dimensionality for our experiments. We use the SAE with 131K dimension from layer 25 for Llama3.1-8B, 65K dimension from layer 25 for Gemma2-2B, and 131K dimension from layer 35 for Gemma2-9B. In our experiments, we choose the top 15 SAE latent features with the highest sensitivity scores for steering, as He et al. (2025b) demonstrate that the model achieves the optimal performance when steering is performed using 15 latent dimensions. Other hyperparameter choices for all the experimental settings are summarized in Table 4 in Appendix C, where we further discuss the influence of different hyperparameters on the steering effect.
@@ 4.2 Benchmark Performance Comparison
Effectiveness of the Proposed SRPS. As shown in Table 1, our steering approach demonstrates superior performance, outperforming the original prompting in 25 out of 27 evaluation settings and achieving comparable results in the remaining two settings on the CSQA dataset. Compared to the prompt-based role-playing baseline, our method achieves better results in 17 out of 27 evaluation settings. These results highlight the effectiveness of our SAE-based steering mechanism across models and tasks.
@@ 4.3 SAE Feature Analysis
To understand how the selected SAE features enhance model reasoning under role-playing, we conduct a semantic analysis of their activations using Neuronpedia Lin (2023). Specifically, we analyze features from Layer 35 of Gemma2-9B for arithmetic reasoning and features from Layer 25 of Gemma2-2B for commonsense reasoning, as shown in Table 2. Figure 3 presents the top-1 ranked feature across all evaluation settings in each domain, showing the corresponding positive and negative output logits. We found the extracted features align well with the expected domain knowledge for each reasoning task, validating the interpretability and task relevance of our steering mechanism. Figure 2 illustrates the impact of our steering method on model outputs. See Appendix D for more examples.
@@
In summary, we found that in both arithmetic and commonsense reasoning tasks there exist features with semantics related to “thinking”. This suggests that one reason for the improved reasoning performance achieved through role-play prompting lies in its ability to encourage step-by-step reasoning.
@@ 4.4 Impact of Feature Ranking in Steering
5 Discussion
In this section, we analyze the instability of role-play prompting and explain why SAE-based role-play steering consistently improves model performance, highlighting its advantages in controllability, stability, and generalizability across model sizes and tasks.
Underlying Reasons Behind the Consistent Effectiveness of Role-Play Steering. While role-play prompting exhibits highly unstable performance across various evaluation settings, our SAE-based steering approach consistently delivers performance gains. The core advantage of our method lies in its controllability. Unlike prompting, where the input text is fixed once applied and cannot be adjusted dynamically, our approach allows fine-grained control over the model’s behavior through tunable parameters. Moreover, prompts often include irrelevant elements such as stop words or punctuation, which may introduce noise and interfere with the model’s reasoning process, resulting in instability or performance degradation. In contrast, our SAE-based steering identifies and applies only those latent features that are strongly correlated with reasoning improvements. By precisely adjusting both the direction and strength of steering, our method amplifies beneficial components while suppressing irrelevant or harmful activations, resulting in more stable and reliable performance improvements.
Another key reason for the superiority of SAE-based steering lies in its broader applicability, particularly for smaller LLMs. These models often lack strong instruction-following capabilities and are not typically fine-tuned on role-play scenarios. As a result, appending complex prompts may exceed the model’s ability to represent information and confuse the model, ultimately harming its performance. In such cases, prompting can disrupt the model’s inherent reasoning abilities rather than enhance them. In contrast, our steering method operates directly on internal representations and does not rely on the model’s ability to interpret natural language instructions. This makes it especially well-suited for smaller models or those without extensive instruction tuning, as it can still induce targeted behavioral changes without relying on external textual guidance. Consequently, the proposed SAE-based steering proves to be a more stable intervention mechanism across different model sizes and capabilities.
@@ 6 Conclusion
In this work, we have introduced an efficient and interpretable steering framework that leverages features extracted from a pretrained SAE to guide LLMs toward role-specific behavior. By constructing steering vectors from highly relevant features and injecting them into the model’s residual stream, our method enables fine-grained control over the model’s internal state without additional training. Extensive experiments across multiple models and reasoning benchmarks demonstrate that our approach consistently enhances model performance, yielding results on par with prompt-based role-playing. In addition, we analyze the semantic meanings of the selected features and reveal that one reason role-playing enhances the reasoning ability of LLMs is its capacity to encourage step-by-step reasoning. Lastly, we evaluate steering performance using $k$ features selected from different ranking positions, and observe a declining trend in accuracy gains as lower-ranked features are used, validating the effectiveness of our feature selection method. Overall, our work demonstrates that steering via SAE enables implicit injection of role-playing information into the model, providing a more stable and interpretable alternative to role-play prompting for improving model reasoning ability.
@@ A.3 Answer Extraction
Appendix B Performance of Role-Play Prompting
@@
Appendix C Hyperparameters of SAE Steering
For each dataset, we perform a grid search over the first 100 test questions to identify relatively optimal parameter combinations. We report all the hyperparameters used for steering across different evaluation settings on the three benchmarks. As the number of features $k$ selected for steering is fixed at 15 in all experiments, we focus on comparing the remaining hyperparameters ($\theta$, $\beta$, and $\lambda$) in Table 4. Specifically, $\theta$ is the activation threshold used to determine whether a feature is considered active; $\beta$ is a weighting coefficient that balances the contribution of activation frequency and strength in feature selection; and $\lambda$ is a scaling factor that controls the strength of the steering vector when injected into the model’s residual stream.
@@ Appendix C Hyperparameters of SAE Steering
Among the three models, Llama3.1-8B is the most sensitive to activation perturbation, and small changes in parameter values can lead to significant differences in steering performance. In contrast, the Gemma-2 family models, particularly Gemma2-9B, is less sensitive to activation modification. Among the three hyperparameters, $\beta$ and $\lambda$ have the greatest impact on steering performance. For models that are more robust to activation intervention, such as Gemma2-9B, larger values of $\beta$ and $\lambda$ (around 10) are typically preferred, and parameter tuning can be performed with larger step sizes (e.g., increments of 5). In contrast, for more steering-sensitive models like Llama3.1-8B, smaller values of $\beta$ and $\lambda$ (around 5) tend to yield better results, and finer-grained tuning with smaller steps (e.g., increments of 1) is recommended. In general, $\theta$ has minimal influence on steering performance and can be set to a default value (e.g., 0.2 or 0.3) to maintain the effectiveness of the steering vectors.